OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#12050 — Lille Optic Loop
Scheduled Maintenance Report for Web Cloud
Completed
Maintenance on the fiber optics is being performed by the operator in Lille, no impact for clients.

Update(s):

Date: 2014-11-18 09:05:34 UTC
The optic loop is secure again.
There will be ongoing maintenance tonight from midnight to 6am, still no impact to customers.
Posted Nov 18, 2014 - 09:03 UTC